Show all work. A aircraft travels 20 miles with a tailwind in 3 hours. It returns in 4 hours. Find the speed of the aircraft in calm wind AND rate of the wind.
Hints: Rate x Times = Distance
3(R+W)=20 or R+W=20/3
4(R-W)=20 or R-W=5

Answer by venugopalramana(3286)

Show all work. A aircraft travels 20 miles with a tailwind in 3 hours. It returns in 4 hours. Find the speed of the aircraft in calm wind AND rate of the wind.
Hints: Rate x Times = Distance
3(R+W)=20 or R+W=20/3..................1
4(R-W)=20 or R-W=5 .........................2
EQN.1+EQN.2..
2R=5+20/3 =35/3
R=35/6.MPH.....SPEED OF AIR CRAFT IN CALM WIND
35/6+W = 20/3
W = 20/3 - 35/6 = 5/6 MPH...SPEED OF WIND
ANY WAY THE FIGURES LOOK ABSURDLY LOW!!FOR AN AIR CRAFT